OpenStudy (kirbykirby):

How to find integral of this abstract function? \(\int_{-\infty}^{u}n(n-1)[F(u)-F(v)]^{n-2}f(u)f(v)dv\)

OpenStudy (kirbykirby):

I got as far as pulling out the constants: \(n(n-1)f(u)\int_{-\infty}^{u}[F(u)-F(v)]^{n-2}f(v)dv\)

OpenStudy (kirbykirby):

And also, F is the antiderivative of f

OpenStudy (kirbykirby):

Apparently the answer is \(nf(u)[F(u)]^{n-1}\)

OpenStudy (abb0t):

\[\int\limits a^n = \frac{ a^{n+1} }{ n+1 }+k\]

OpenStudy (abb0t):

\[n-2+1 = n-1\]

OpenStudy (abb0t):

once you have your answer, evaluate from 0 to u and then take the limit as it approaches ∞

OpenStudy (kirbykirby):

I'm not sure how to deal with the f(v) though :S

OpenStudy (kirbykirby):

since it's being multiplied

OpenStudy (anonymous):

\[n(n-1)\int_{-\infty}^u\left[F(u)-F(v)\right]^{n-2}f(u)f(v)\;dv\] Assuming F(u) and F(v) are antiderivatives of f(u) and f(v), respectively, make the substitution \[t=F(u)-F(v)\\ dt=-f(v)\;dv\] That should help you get the form @abb0t was suggesting.

OpenStudy (kirbykirby):

:o Oh I see what you mean.. that's a good idea :)